# Superconducting qubits at the utility scale: the potential and limitations of modularity

### v2

@@ I.2 High-level macro-architecture for modular graph state processing
We consider a physical layout for FTQC consisting of sets of rotated surface code patches, which are tiled together to create all intra-module components (e.g., a group of patches becomes a graph-state memory register, and others become an MSD factory). Each patch contains $2d^{2}$ physical qubits. While the rotated surface code only requires $12d^{2}-$ 1 physical qubits [12, 16, 17, 18], we adapt a physical size of $2d^{2}$ for all patches to streamline our scaling calculations.
@@
Executing the quantum algorithm based on the ASG formalism and time-direction widgetization inherently results in a macro-architecture that divides the hardware into sets of modules. The question is how to lay out modules and assign them to potentially interleaving operations. During a steady state, the execution of the quantum algorithm comprises three main steps: 1) preparation or creation of the graph, which initializes a quantum state based on the original entangling operations, 2) consumption of the graph, which executes the original non-Cliffords, and, 3) following every consumption step, the handover of the graph through Bell-pair teleportations or stitching of output nodes from the preceding widget to the input nodes of the current widget.
@@
Figure 2: The high-level graph state processing cycle: the algorithm is divided into time-sliced widgets interleaved across two module sets in the time direction. Modules are arranged over two legs, A and B, of a ladder structure (in this example, each leg of the ladder contains three modules). Each module set manages graph preparation, consumption, and the local magic state distillation. Light purples mean the module distills, routes, or consumes $T$-states. The algorithm interconnects module sets A and B through the teleportation of output nodes from previous widgets in one leg to the input nodes of the next ones in the other leg.
@@
Figure 3: The high-level macro-architecture consists of two sets of modules arranged like the legs of a ladder structure. Following the example in fig. 2, each leg of the ladder contains three modules. Portions of the algorithm are segmented and executed solely on individual modules along the same leg, where intra-modular interactions, characterized by high connectivity and fast gate times, dominate computations. All modules communicate through coherent links at a reduced rate and with lower connectivity, exclusively for graph preparation and handover.
@@
Step 2 for the current widget and Step 1 for the next can co-occur in the graph state processing. This interleaved execution structure enables us to explore a degree of parallelism in a distributed FTQC. We propose distributing the interleaving preparation and consumption steps into two physically distinct sets of modules positioned along the two legs of a ladder structure. The graph processing cycle is streamlined because each set of modules alternates between only preparing or consuming the graph at each step. In contrast, the other sets of modules do the opposite on the adjacent widget, as illustrated in fig. 2. During all preparation and consumption steps, the FTQC performs MSD within the modules and fills in a dedicated $T$-transfer bus as required. The timescales for $T$-state distillation and inter-module operations are typically one order of magnitude worse than other surface code operations, including graph preparations, so MSD will remain local to improve the speed of computations. Only the consumption stage requires the FTQC to apply the $T$ states; since the graph preparations and MSD are performed independently, we need to account for possible preparation, MSD, and other delays in the consumption stage of each widget. For the consumption steps, some graph-state nodes need a single $T$-basis measurement, while others require multiple $T$-basis measurements to execute non-Clifford rotations in the original algorithm. Subsequently, FTQC teleports the current graph’s output nodes to the newly prepared widget’s input nodes. This cycle continues until FTQC processes all widgets in the complete algorithm (see also section I.4.6).
@@
The architecture must be scalable for algorithms requiring more space-time resources than a single module or time step. Furthermore, MSD is arguably the most resource-intensive component of a universal FTQC, and we set an equal number of MSD factories supplying the $T$ states locally in each module. To address both issues, we propose a high-level ladder-style macro-architecture illustrated in fig. 3 (see fig. 6 for the intra-module layout). The quantum memory register will be shared equally among all modules on the same leg of the ladder at any given time. Here, we can utilize the faster native connectivity of a lattice of superconducting qubits within a module for graph consumption and MSD, while reduced connectivity or slower interactions are employed for some graph preparation operations and teleportation of the output nodes of graph states on one leg of the ladder to the input nodes on the other leg.
We can estimate the time required to execute the complete schedule of operations across the system by considering all local and inter-module operations. To account for the inter-module operations, we abstract away the characteristics of the teleportation interaction between modules as a single time referred to as intermodule handover time and set it to $t_{\text{inter}}=1~{}\mu$s. We incorporate this characteristic time to measure the cost of scheduled operations crossing a module boundary, estimating the total fault-tolerant handover time.
I.3 Superconducting hardware micro-architecture and lower-level components
@@
We propose a micro-architecture concept (see fig. 6 for an illustrative example) that informs a system-level model, which must, at a minimum, include the following components:
@@
    A lattice of rotated surface code qubits used as logical patches for building all components,
@@
    Continuous buses of logical ancilla qubits that couples with, connects to, and operate on memory qubits while routing $T$-states for graph consumption.
@@
    A continuous bus of ancillary qubits for queuing $T$-states must maintain a broad enough connection to the preceding ancillary qubits to teleport $T$-states to them on demand.

### v3

@@ II.2 High-level macro-architecture for modular graph state processing
We consider a macro-architectural model that, at its core, uses superconducting qubits on square lattices and the logical layout shown in fig.˜1(bottom). This enables us to estimate the resources required to execute the FT algorithm and understand the associated trade-offs. To this end, we consider a physical layout for FTQC consisting of sets of rotated surface code patches, which are tiled together to create all intra-module components. For example, a group of patches forms a graph-state memory register, while others are designated for the $T$-transfer bus. We assume that each patch contains $2d^{2}$ physical qubits. Although the rotated surface code requires only $2d^{2}-1$ physical qubits [39, 52, 18, 58], we adopt a physical size of $2d^{2}$ for all patches to streamline our approximate scaling calculations, ignoring non-dominant spatial terms and tiling details.
@@
Executing the quantum algorithm based on the ASG formalism and time-direction widgetization inherently yields a macro-architecture that partitions the hardware into sets of modules that execute widgets in an interleaved fashion. The question is how to lay out modules and assign them to potentially interleaving operations. In steady state, the execution of the quantum algorithm comprises three main steps: 1) preparation or creation of the graph, which initializes a quantum state using the original nontrivial Clifford and entangling operations, 2) consumption of the graph, which executes the original non-Cliffords, and 3) following every consumption step, the handover of the graph through Bell-pair teleportations or by stitching output nodes from the preceding widget to the input nodes of the current widget.
@@
Figure 2: The high-level graph state processing cycle: the algorithm is divided into time-sliced widgets that are interleaved across two module sets in the time direction. Each set, consisting of three modules in this example, interleaves graph-state preparation with graph-state consumption. Graph-state preparation consists solely of Clifford operations (blue), while consumption utilizes $T$-state resources (purple). Teleportation moves logical state between module sets, from the output of one graph-state consumption to the input of the next. Teleportation is also used to facilitate operations that cross module boundaries within the set.
@@
Figure 3: The high-level macro-architecture consists of two sets of 1-million physical qubit (1MQ) modules arranged like the legs of a ladder. Following the example in fig.˜2, each leg contains three modules, each following the layout of fig.˜1(bottom). Portions of the algorithm are divided and executed solely within modules on the same leg, where intra-module interactions with high connectivity and short gate times dominate. All modules communicate via width-$d$ “coherent interconnects” (or links) at a lower rate and with reduced connectivity, used only for graph preparation (vertical interconnects) and state handover (horizontal interconnects).
@@
Step 2 for the current widget and Step 1 for the widget can occur concurrently during graph state processing. This interleaved execution structure enables us to explore a degree of parallelism in a distributed FTQC. We propose distributing the interleaving preparation and consumption steps across two physically distinct sets of modules positioned along the two legs of a ladder structure. Teleportation may be used to mediate graph processing operations spanning multiple modules, but these should be minimized in scheduling. The graph processing cycle is streamlined because each set of modules alternates between only preparing or consuming the graph at each step. In contrast, the other sets of modules do the opposite on the adjacent widget, as illustrated in fig.˜2. During all preparation and consumption steps, the FTQC performs MSD within the modules and fills a dedicated $T$-transfer bus as required. The timescales for $T$-state distillation and inter-module operations are typically one order of magnitude worse than those for other surface code operations, including graph preparations, so MSD remains local to individual modules to improve computation speed. Only the consumption stage requires the FTQC to apply the $T$ states; since graph preparations and MSD are performed independently, we need to account for possible preparation, MSD, and other delays in each widget’s consumption stage. For the consumption steps, some graph-state nodes require a single $T$-basis measurement, while others require multiple $T$-basis measurements to execute non-Clifford rotations in the original algorithm. Subsequently, the FTQC teleports the current graph’s output nodes to the newly prepared widget’s input nodes. This cycle continues until the FTQC processes all widgets in the complete algorithm (see also section˜II.4.7).
@@
The architecture must be scalable to accommodate algorithms that require more space-time resources than a single module or time step. Furthermore, MSD is arguably the most resource-intensive component of a universal FTQC, and we assume that each module has an equal number of $T$-factories supplying $T$-states locally. To address both issues, we propose a ladder-style macro-architecture illustrated in fig.˜3 (see fig.˜1 for the intra-module layout). The quantum memory register will be shared equally among all modules on the same ladder leg at any given time. Here, we can use the faster native connectivity of a lattice of superconducting qubits within a module for graph consumption and MSD, while reduced connectivity or slower interactions are used for some graph preparation operations and for teleporting the output nodes of graph states on one leg of the ladder to the input nodes on the other leg.
[…]
II.3 Superconducting hardware micro-architecture and lower-level components
@@
We propose a logical micro-architecture concept (see fig.˜1) that informs a system-level model, which, at a minimum, must include the following components:
  1. 0.
    At its core, a lattice of rotated surface-code physical qubits, used as logical patches to build the following distinct subsystems.
@@
    Continuous buses of logical auxiliary qubits that couple with, connect to, and operate on logical memory qubits while routing $T$-states for graph consumption.
@@
    Another continuous bus of logical auxiliary qubits for queuing and moving $T$-states must maintain a broad enough connection to the preceding auxiliary qubits to transfer $T$-states to them on demand.
@@
    $T$-factories.